OpenStudy (anonymous):

Write an equation of the line, in point-slope form, that passes through the two points.

OpenStudy (hba):

@Falkqwer Lemme provide you the formula \[\huge\ (y-y1)=\frac{ y_2-y_1 }{ x_2-x_1 }(x-x_1)\] (x1,y1)= (–13, 9), (x2,y2)= (11, –3)

OpenStudy (anonymous):

Its b :) am I right?

OpenStudy (hba):

Yes

OpenStudy (hba):

@Falkqwer Don't worry it is correct.

OpenStudy (anonymous):

OK thank you :)

OpenStudy (anonymous):

I deleted them after he said I was wrong/ right....

OpenStudy (hba):

She put it before in the question. See under my formula Now i guess she removed them from the ques

OpenStudy (hba):

@Hero Stop calling me psychic -_-

OpenStudy (anonymous):

That was not funny...I ... :'(

OpenStudy (hba):

@Hero Shu.